Please help me with my problem.
The product of two consecutive integers is 272. Find the integers. Show all work necessary to arrive to your answer. Please show work.

Answer by stanbon(75887)

The product of two consecutive integers is 272. Find the integers. Show all work necessary to arrive to your answer. Please show work.
---------------------------
1st: x
2nd: x+1
-----------
Equatiion:
x(x+1) = 272
x^2+x-272 = 0
(x-16)(x+17) = 0
---
If x = 16
Then x+1 = 17
----
If x = -17
The x+1 = -16
==================
Cheers,
Stan H.
===================